# DWaste: Greener AI for Waste Sorting using Mobile and Edge Devices

### v1

@@ Abstract
The rise of convenience packaging has led to generation of enormous waste, making efficient waste sorting crucial for sustainable waste management. To address this, we developed DWaste, a computer vision-powered platform designed for real-time waste sorting on resource-constrained smartphones and edge devices, including offline functionality. We benchmarked various image classification models (EfficientNetV2S/M, ResNet50/101, MobileNet) and object detection (YOLOv8n, YOLOv11n) using a subset of our own waste data set and annotated it using the custom tool Annotated Lab. We found a clear trade-off between accuracy and resource consumption: the best classifier, EfficientNetV2S, achieved high accuracy ($\approx 96\%$) but suffered from high latency ($\approx 0.22\text{s}$) and elevated carbon emissions. In contrast, lightweight object detection models delivered strong performance (up to 77% $7\%\text{ mAP}$) with ultra-fast inference ($\approx 0.03\text{s}$) and significantly smaller model sizes ($7\text{MB}$), making them ideal for real-time, low-power use. Model quantization further maximized efficiency, substantially reducing model size and VRAM usage by up to $75\%$. Our work demonstrates the successful implementation of "Greener AI" models to support real-time, sustainable waste sorting on edge devices.
@@ 2.1 Dataset and Preprocessing
In this study, we used our garbage dataset [16] focusing on a subset of seven categories deemed critical for recycling efficiency: biological, cardboard, glass, metal, paper, plastic, and trash. These images were collected from the internet, DWaste platform, and community submissions. All images were annotated with category labels and bounding boxes using Annotate Lab [17]. The final processed dataset consists of 11,163 images and 19,700 bounding box instances shown below Figure 1.
@@ 2.2 Model Training
We evaluated both classification (EfficientNetV2S/M, MobileNet, ResNet50/101) and object detection (YOLOv8n, YOLOv11n) architectures using a transfer learning approach. All classification models were initialized with weight pre-trained on the ImageNet dataset. All models were trained for 20 epochs using NVIDIA Tesla T4x2 GPU in Kaggle. The performance of each model was benchmarked using standard performance metrics including Accuracy, Precision, Recall, F1-score, and Mean Average Precision (mAP). With our focus on sustainable AI deployment, we conducted detailed analysis of VRAM usage, model size, and carbon emissions across various phases of the workflow. Carbon emissions were monitored using CodeCarbon library [20]. The resulting full-precision models were further optimized via quantization, a technique known to achieve up to 95% reduction in the number of parameters and model size [21], thereby significantly lowering energy use. Figure 3 depicts the overall process, while the corresponding experiments are available on Github [^1] https://github.com/sumn2u/AIMS-2025.
Figure 3: End-to-end machine learning pipeline
@@ 3 Results and Discussion
YOLOv8n 75.50 0.78 0.65 0.75 0.76 6.5 3.1
YOLOv11n 77.10 0.77 0.69 0.77 0.77 5.4 2.8
@@
The high-performance classification models EfficientNetV2M and EfficientNetV2S showed highest accuracy ($\approx 95-96\%$) but was constrained by larger model sizes (216MB and 84.3MB respectively) and consequently resulted in higher training and deployment emissions shown in Figure 4. Similarly, ResNet101 and ResNet50 delivered strong accuracy (91-92%) but were also penalized by substantial size and higher emissions. In contrast, MobileNet prioritized resource efficiency, exhibiting the smallest initial size (14.7 MB, reduced to 3.5 MB after quantization) with minimal energy usage and lowest accuracy ($\approx 67\%$). The lightweight object detection models, YOLOv8n and YOLOv11n demonstrated the most balanced trade-off, achieving accuracy $75\text{-}77\%$ accuracy with impressive quantized sizes ($\approx 3$MB).
@@
Notably, YOLOv11n emerged as the optimal architecture for edge deployment, attaining the best mean Average Precision (mAP = 0.77) with the smallest quantized footprint (2.8 MB) and the fastest speed, confirming its efficacy for resource-constrained applications with minimal carbon emissions.
This optimized model has been successfully deployed to both the DWaste mobile application and a dedicated edge device for a real time waste object. An example of waste detection using both the app and edge device is shown in Figure 8.
@@ 4 Conclusion
This study successfully quantified the inherent trade-off between model accuracy and energy efficiency for deep learning-based waste sorting systems. Our results clearly demonstrate that while larger architectures (EfficientNetV2S/M, ResNet101/50) achieve superior classification accuracy, they demand greater computational resources and incur significantly higher carbon emissions. Conversely, lightweight object detection models, specifically YOLOv11n and YOLOv8n, strike a crucial balance, offering strong real-time performance (mAP 77%) with minimal resource overhead. Furthermore, we confirmed that the application of model quantization dramatically aids deployment, consistently and substantially reducing VRAM usage and model size, thus directly lowering the energy required for inference on edge and mobile devices. Given these findings, the YOLOv11n model, which achieved the best combination of accuracy, inference speed, and minimal resource usage (smallest quantized footprint), was selected and successfully embedded into the DWaste mobile app and a dedicated edge device for practical waste sorting implementation. Future work should focus on refining lightweight model accuracy with an expanded dataset and conducting a longitudinal study to evaluate the long-term effectiveness and economic impact of the deployed app and the edge system.

### v2

@@ Abstract
The rise of convenience packaging has led to generation of enormous waste, making efficient waste sorting crucial for sustainable waste management. To address this, we developed DWaste, a computer vision-powered platform designed for real-time waste sorting on resource-constrained smartphones and edge devices, including offline functionality. We benchmarked various image classification models (EfficientNetV2S/M, ResNet50/101, MobileNet) and object detection (YOLOv8n, YOLOv11n) including our purposed YOLOv8n-CBAM model using our annotated dataset designed for recycling. We found a clear trade-off between accuracy and resource consumption: the best classifier, EfficientNetV2S, achieved high accuracy ($\approx 96\%$) but suffered from high latency ($\approx 0.22\text{s}$) and elevated carbon emissions. In contrast, lightweight object detection models delivered strong performance (up to $80\%\text{ mAP}$) with ultra-fast inference ($\approx 0.03\text{s}$) and significantly smaller model sizes ($7\text{MB}$), making them ideal for real-time, low-power use. Model quantization further maximized efficiency, substantially reducing model size and VRAM usage by up to $75\%$. Our work demonstrates the successful implementation of "Greener AI" models to support real-time, sustainable waste sorting on edge devices.
@@ 1 Introduction
While our own past study focused on benchmarking models based on accuracy and carbon emission to determine a "greener" classification model, it was limited to classification tasks only [16]. This paper aims to address these challenges by systematically addressing the trade-off in deploying advanced DL models. We benchmark state of the art classification and object detection models including our own model that uses CBAM inhanced backbone and use model quantization as the primary optimization technique, precisely benchmarking the reduction in VRAM usage, model size, and carbon emission to validate the path toward a "Greener AI" solution for waste sorting. Later, the greener model is deployed to mobile apps and edge devices.
@@ 2.1 Dataset and Preprocessing
In this study, we used our garbage dataset [17] focusing on seven categories deemed critical for recycling efficiency: biological, cardboard, glass, metal, paper, plastic, and trash. These images were collected from the internet, DWaste platform, and community submissions. All images were annotated with category labels and bounding boxes using Annotate Lab [18]. The final processed dataset consists of 11,163 images and 19,700 bounding box instances shown below Figure 1.
@@ 2.2 Model Training
We evaluated both classification (EfficientNetV2S/M, MobileNet, ResNet50/101) and object detection (YOLOv8n, YOLOv11n) architectures including our proposed YOLOv8n-CBAM model shown in Figure 3, using a transfer learning approach.
InputImageYOLOv8nBackboneCBAM(Channel + Spatial Attention)DetectionHeadOutput
Figure 3: Architecture of the proposed Improved YOLOv8n model with CBAM-enhanced backbone
All classification models were initialized with weight pre-trained on the ImageNet dataset. All models were trained for 20 epochs using NVIDIA Tesla T4x2 GPU in Kaggle. The performance of each model was benchmarked using standard performance metrics including Accuracy, Precision, Recall, F1-score, and Mean Average Precision (mAP). With our focus on sustainable AI deployment, we conducted detailed analysis of VRAM usage, model size, and carbon emissions across various phases of the workflow. Carbon emissions were monitored using CodeCarbon library [21]. The resulting full-precision models were further optimized via quantization, a technique known to achieve up to 95% reduction in the number of parameters and model size [22], thereby significantly lowering energy use.
@@ 3 Results and Discussion
YOLOv8n - 0.78 0.65 0.75 0.76 6.5 3.1
YOLOv11n - 0.77 0.69 0.77 0.77 5.4 2.8
YOLOv8n-CBAM - 0.78 0.73 0.80 0.80 6.1 3.5
@@
The high-performance classification models EfficientNetV2M and EfficientNetV2S showed highest accuracy ($\approx 95-96\%$) but was constrained by larger model sizes (216MB and 84.3MB respectively) and consequently resulted in higher training and deployment emissions shown in Figure 4. Similarly, ResNet101 and ResNet50 delivered strong accuracy (91-92%) but were also penalized by substantial size and higher emissions. In contrast, MobileNet prioritized resource efficiency, exhibiting the smallest initial size (14.7 MB, reduced to 3.5 MB after quantization) with minimal energy usage and lowest accuracy ($\approx 67\%$). The lightweight object detection models, YOLOv8n-CBAM and YOLOv11n demonstrated the most balanced trade-off, achieving $78\%$ precision with impressive quantized sizes ( $3.6,\text{MB}$).
@@
Notably, YOLOv11n emerged as the optimal architecture for edge deployment, attaining the best mean Average Precision (mAP = 0.77) with the smallest quantized footprint (2.8 MB) and the fastest speed, confirming its efficacy for resource-constrained applications with minimal carbon emissions. Our proposed model performed best in terms of precision, recall, F1-score, and mAP before quantization, and also had low VRAM usage with compared to YOLOv8n and YOLOv11n. However, its VRAM usage increased significantly after quantization.
The YOLOv8n-CBAM model has been successfully deployed to both the DWaste mobile application and a dedicated edge device for a real time waste object. An example of waste detection using both the app and edge device is shown in Figure 8.
@@ 4 Conclusion
This study successfully quantified the inherent trade-off between model accuracy and energy efficiency for deep learning-based waste sorting systems. Our results clearly demonstrate that while larger architectures (EfficientNetV2S/M, ResNet101/50) achieve superior classification accuracy, they demand greater computational resources and incur significantly higher carbon emissions. Conversely, lightweight object detection models, specifically YOLOv11n and YOLOv8n-CBAM, strike a crucial balance, offering strong real-time performance (mAP 77% and 78%) with minimal resource overhead. Furthermore, we confirmed that the application of model quantization dramatically aids deployment, consistently and substantially reducing VRAM usage and model size, thus directly lowering the energy required for inference on edge and mobile devices. Given these findings, the YOLOv8n-CBAM model, which achieved the best combination of accuracy, inference speed, and minimal resource usage, was selected and successfully embedded into the DWaste mobile app and a dedicated edge device for practical waste sorting implementation. Future work should focus on refining lightweight model accuracy with an expanded dataset, and conducting a longitudinal study to evaluate the long-term effectiveness and economic impact of the deployed app and the edge system.